OpenStudy (anonymous):

Investments. A self-employed contractor nearing retirement made two investments totaling $15,000. In one year, these investments yielded $1023 in simple interest. Part of the money was invested at 6% and the rest at 7.5%. How much was invested at each rate?

OpenStudy (saifoo.khan):

Do you know the simple interest formula?

OpenStudy (anonymous):

I do not remember...

OpenStudy (anonymous):

I really appreciate your help

OpenStudy (anonymous):

So, using that formula, how would you set up the equation?

OpenStudy (saifoo.khan):

Ah. wait. this is a bit different.

OpenStudy (saifoo.khan):

So: Let's suppose x and y is the variable amount. x + y = 15000 And 0.06x + 0.075y = 1023

OpenStudy (anonymous):

Ok...

OpenStudy (saifoo.khan):

Can you solve further/

OpenStudy (anonymous):

So, x+y=15,000 0.06x+0.075y=1023 x=-y+15000 0.06(-y+15000)+0.075=1023 -0.06y+900+0.075y=1023 -0.015y+900=1023 -0.015y=-900+1023 -0.015y=123 -0.015y/-0.015=123/-0.015 y=-8200 y=-(-8200) y=8200 x=-8200+15000 x=6800

OpenStudy (anonymous):

Is this correct?

OpenStudy (saifoo.khan):

Perfect.

OpenStudy (anonymous):

Thank you!!!! You are awesome!!!!